OpenStudy (anonymous):

the sum of two consecutive integers is -49. write an equation that models this situation and find the values of the two integers.

OpenStudy (kirbykirby):

Say \(n\) is an integer. Then you can say that the integer right after \(n\) is \(n+1\) (If you're not sure why: Say you take "4" as n, so n=4, then the integer after 4 is "5", which is "n+1=4+1") So now you want \(n+(n+1)=-49\)

OpenStudy (kirbykirby):

\(n+n+1=-49\implies 2n+1=-49 \implies n=-25\)